Find the first and second derivatives of the following and simplify whenever possible:



x = t ^ 2 * e ^ t y = t In t

1
Expert's answer
2021-11-29T13:15:56-0500
"\begin{matrix}
 x=t^2e^t \\
 y=t\ln t
\end{matrix}"

"\dfrac{dx}{dt}=2te^t+t^2e^t"

"\dfrac{dy}{dt}=\ln t+t(\dfrac{1}{t})=\ln t+1"

"\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{\ln t+1}{2te^t+t^2e^t}"

"\dfrac{d}{dt}(\dfrac{dy}{dx})=\dfrac{\dfrac{1}{t}(2te^t+t^2e^t)-(2e^t+4te^t+t^2e^t)(\ln t+1)}{(2te^t+t^2e^t)^2}"

"=\dfrac{2e^t+te^t-(2e^t+4te^t+t^2e^t)\ln t-2e^t-4te^t-t^2e^t}{(2te^t+t^2e^t)^2}"

"=-\dfrac{3te^t+t^2e^t+(2e^t+4te^t+t^2e^t)\ln t}{(2te^t+t^2e^t)^2}"

"\dfrac{d^2y}{dx^2}=\dfrac{\dfrac{d}{dt}(\dfrac{dy}{dx})}{dx/dt}"

"=-\dfrac{3te^t+t^2e^t+(2e^t+4te^t+t^2e^t)\ln t}{(2te^t+t^2e^t)^3}"